OpenStudy (anonymous):

how to solve the system {y=x^2+4, xy=6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

y= 6/x plug into the first eqn 6/x = x^2 + 4 6 = x^3 + 4x

OpenStudy (anonymous):

Then what???

OpenStudy (anonymous):

then solve the cubic eqn